The probability of an event A is 3-4x. Find the value of x which can be and cannot be.
Answer by Theo(13342)

the probability of an event has to be greater than or equal to 0 and less than or equal to 1.

therefore:

3-4x >= 0 and 3-4x <= 1

when 3-4x >= 0, solve for x to get x <= 3/4.

when 3-4x <= 1, solve for x to get x >= 1/2.

x has to be greater than or equal to 1/2 and less than or equal to 3/4.

x cannot be less than 1/2 and cannot be greater than 3/4.

when x = 1/2, 3-4x becomes 3-4*1/2 which becomes 3-2 which becomes 1.

when x = 3/4, 3-4x becomes 3-4*3/4 which becomes 3-3 which becomes 0.

if x is less than 1/2, say 0, then you get 3-4*0 = 3-0 = 3 which is invalid.

if x is greater than 3/4, say 1, then you get 3-4*1 = -1 which is invalid.

when x is between 1/2 and 3/4, say 5/8, then you get 3-4*5/8 = 3-20/8 = (24-20)/8 = 4/8 = 1/2 which is between 0 and 1, therefore valid.